OpenStudy (anonymous):

Is △TVS scalene, isosceles, or equilateral? The vertices are T(1, 1), V(4, 0), and S(3, 5).

OpenStudy (anonymous):

TV = sqrt(3^2 + 1) = aqrt10 VS = sqrt(1 + 5^2) = sqrt26 ST = sqrt(2^2 + 4^2) = sqrt20 3 sides of diffrent lengths so the triangle is scalene

OpenStudy (anonymous):

Thank You!

OpenStudy (anonymous):

yw